OpenStudy (anonymous):

How do you graph the function f(x)=6^x? Explain it to me, please.

OpenStudy (anonymous):

You use a graphing calculator.

OpenStudy (anonymous):

Is there any way to do it without one?

OpenStudy (richyw):

are you trying to sketch it, or do you want to make a proper graph of the function?

OpenStudy (ranga):

a^x where a is a positive number greater than 1 all have the similar plot as e^x. The function will always be positive and so the plot will be above the x-axis. It will cross the y-axis at (0,1) When x-> -infinity, f(x) -> 0.

OpenStudy (anonymous):

I'm trying to sketch it but I don't have any graphing technology

OpenStudy (ranga):



OpenStudy (richyw):

if you just want to make a sketch of it, then what ranga said is good. This tells you all the important stuff about the function. of course you could always make a table of values, and plot the points.

OpenStudy (the_fizicx99):

Just make a table, for f(x), 1,2,3,4 etc. Plot them on the graph: f(1) = 6^1 = 6 (1,6) f(2) = 6^2 = 36 (2,36) ... and so on

OpenStudy (richyw):

also, assuming you are on a computer, you do have graphing technology available.

OpenStudy (anonymous):

It's a long story but I don't have it. But thank all of you guys for helping!! I think I can do it on my own now! :)

OpenStudy (the_fizicx99):

Have fun :p

OpenStudy (anonymous):

Oh, I just love math.. ;D